OpenStudy (anonymous):

What is the solution set of the given system? 2x - y - 4 = 0 3x + y - 9 = 0 Choices;{(6/5, 13/5)} {(13/5, 6/5)} {(5, 6)} {(6, 5)}

OpenStudy (anonymous):

2x - y - 4 = 0 y = 2x - 4 Substituting this in second equation: 3x + 2x - 4 - 9 = 0 5x = 13 x = 13/5. So its second one.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you are welcomed.